MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2006/11/16
OBSERVER: Kamiya
CONDITION: cloudy
COMMENT: I adjusted (removed and inserted) the video board for chip2, because chip2 was noisy. After that chip2 had no strange noise, but bias level became low (about 150ADU). So I adjusted bias level (test118.lod, old is test115.lod). D1814(0sec) is OK. I take flats(F1303-1312) to estimate CCD gain.
Compressor gas pressure: MPa
Expected Observation Time: START 22:17 END 04:46
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | - | 0 | 0 | 0 |
LMC | - | 0 | 0 | 0 |
SMC | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 0 | 0 | 0 (00h00m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 |
0 | 0 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CHECK SHEET
Time | Status | Weather |
22:00 | waiting | cloudy |
23:00 | waiting | cloudy |
0:00 | waiting | cloudy |
1:00 | waiting | cloudy |
2:00 | waiting | cloudy, slightly rain |
3:00 | waiting | cloudy |
4:00 | waiting | cloudy |
5:00 | waiting | dawn |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day